1 Determine k for which line (k + 2)x + (k^2 -9)y + 3k^2 - 8k + 5 =0 is 1, parallel to y axis 2, parallel to x axis 3, pass through origin 4, pass through (1,1)

1
Expert's answer
2021-12-13T16:40:02-0500

1.

parallel to "x=0"

then:

"k^2-9=0"

"k=\pm 3"


2.

parallel to "y=0"

then:

"k+2=0"

"k=-2"


3.

"3k^2 - 8k + 5 =0"


"k=\frac{8\pm \sqrt{64-60}}{6}"


"k_1=1,k_2=5\/3"


4.

"(k + 2) + (k^2 -9) + 3k^2 - 8k + 5 =0"

"4k^2-7k-2=0"


"k=\frac{7\pm \sqrt{49+32}}{8}"


"k_1=-1\/4,k_2=2"